OpenStudy (anonymous):

one number is 2 more than another. The difference between their squares is 28. What are the numbers?

OpenStudy (anonymous):

Solve the following for a and b.\[\left\{a+2=b,b^2-a^2=28\right\} \]\[\{a\to 6,b\to 8\} \]